Question 6 a) As a promotional offer, a new gym is offering a 30% discount to those who sign up in the first month. Patty and her 2 siblings take advantage of the offer and sign up for total amount of £210. What were the total savings made? Show your working.

b) Work out the average savings per person.

Question 7 Work out the following: a) 3/4 - 7/9 + 2/3 (4 marks) b) Which is the largest of the following numbers? 0.1, 0.02, 0.003, 0.0004, 0.00005

Question 8 90 men and 60 women were asked if they had watched the latest ‘Expendables’ movie. Altogether 3/5 of the people said Yes. 3/10 of the women said Yes. What percentage of the men said No? (Show your working)

working. 


1
Expert's answer
2021-10-21T14:38:33-0400

6.

a

amount = 210

savings = 30%


Total savings

Total savings = savings * amount

"=30\%\times 210=63"


b

Average savings

"=\frac{63}{3}=21"


7.

a

"\frac{3}{4} - \frac{7}{9} + \frac{2}{3}"


"\frac{3}{4}+\frac{2}{3}=\frac{3\times3+4\times2}{12}=\frac{17}{12}\\"


"\frac{17}{12}-\frac{7}{9}=\frac{3\times17-4\times7}{36}=\frac{23}{36}"


b


"0.1\times 100000=10,000\\ 0.02\times100000=2000\\0.003\times100000=300\\0.0004\times100000=40\\ 0.00005\times100000=5"


The greatest number is 0.1


8

"yes=\frac{3}{5}\times150=90"

women who said yes "=\frac{3}{10}\times60=18"

men who said yes "=90-18=72"

men who said no "=90-72=18"